OpenStudy (anonymous):

Identify the solution of the compound inequality −3y < 12 and y − 2 ≤ 4.

OpenStudy (anonymous):

\[-3y<12 \] solve this y divide by (-3) (this will change direct) result is :-\[y>4\]in the other side if we move 2 to right side : \[y-2\le 4 \] after moving \[y\le 6 \] intersection between them you will get :- \[4<y\le 6 \]

OpenStudy (anonymous):